The coordinates of the midpoint of line segement AB are (-2,4).if the coordinates of point A are (7,10),find the coordinates of point B.
1
Expert's answer
2012-09-27T11:40:22-0400
Let x and y be the coordinates of point B. Then we have the following system of equations:

& (7+x)/2 = -2 (1)
(10+y)/2 =& 4 (2)

Then

(1) ==> x = -2*2 -& 7 = -11;
(2) ==> y =& 4*2 - 10 = -2;

So, the coordinates of point B are (-11,-2).